dot and her cousin have a combined age of 24 years. 3years ago, dot was exactly twice as old as her cousin. how old is dot?
Answer by stanbon(75887)

dot and her cousin have a combined age of 24 years. 3 years ago, dot was exactly twice as old as her cousin. how old is dot?
------------------------------------------------
Equations:
d + c = 24
d-3 = 2(c-3)
----------------
Substitute for d and solve for "c":
24-c - 3 = 2c - 6
3c = 27
c = 9 yrs (cousin's age)
---------------------------
Solve for "d":
d + 9 = 24
d = 15 yrs (dot's age)
===========================
Cheers,
Stan H.
===========